OpenStudy (anonymous):

How is x/(sqrt(x^2+1)) = 1-(1/(sqrt(x^2+1)) ?

OpenStudy (anonymous):

square top and bottom of x/(sqrt(x^2+1) add 1/(1/sqrt(x^2+1) to both sides. they are equivalent

OpenStudy (phi):

Are you supposed to solve for x?

OpenStudy (anonymous):

no I am just trying to figure out how 1-(1/sqrt(x^2+1)) is an alternate way to write x/(sqrt(x^2+1))

OpenStudy (phi):

if I understand this, then you are saying, with x=1, 1/sqrt(2) = 1 - 1/sqrt(2) ?

OpenStudy (anonymous):

yeah, it's part of a limit that i am evaluating as x\[\rightarrow \infty\]

OpenStudy (phi):

What you posted is only true for x=0. Post the original limit problem.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}x/\sqrt{x^2+1}\]

OpenStudy (phi):

It's too long to type. See http://www.wolframalpha.com/input/?i=limit+x-%3E+infinity+x%2Fsqrt%28x%5E2%2B1%29 and type show steps in the upper right corner of the first box.

OpenStudy (anonymous):

Thanks, for your help.

OpenStudy (phi):

*and click show steps in the upper right corner of the first box.

OpenStudy (anonymous):

Yeah i checked it, I'll have to brush up on the power rules for limits :) Thanks for the reply, very helpful.